OpenStudy (anonymous):

multiply square root of 8 by square root of 10 in radical form

OpenStudy (anonymous):

8.94427

OpenStudy (owlcoffee):

Let's first look at the properties of the "radicals". We will call a "radical" to any number or mathematical expression that is inside a square root. So, let's say we have two radicals: \[Q = \sqrt[n]{a}\] \[P=\sqrt[n]{b}\] So, in other words, I want to know what the value of the product of Q and P are: \[Q.P=(\sqrt[n]{a})(\sqrt[n]{b})\] using some exponential properties we get: \[(Q.P)^n = a.b\] Taking the n root on both sides: \[Q.P=\sqrt[n]{a.b}\] And replacing Q and P values: \[\sqrt[n]{a}.\sqrt[n]{b}=\sqrt[n]{b.a}\] So, what we just found is that, any product of radicals, is the same as the radical of the product of the expressions inside of them, cool huh? For example: \[\sqrt{5}.\sqrt{40}=\sqrt{5.40}=\sqrt{200}\]